Please answer those two questions. Thanks.
During Heaton Companys first two yoars of operations, the company reported absorption costing not operating income as follows: S 930,000 1,530,000 527,000 867,000 403,000663,000 Sales (@ $60 per unit) Cost of goods sold (@ $34 per unit) Gross margin Selling and administrative expenses Net operating income 300,700 330,700 $ 102,300 332,300 ·$3 per unit variable; S254.200 fixed each year. The companys $34 unit product cost is computed as follows: Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead ($369,000 20,500 units)18 $ 34 Absorption costing unit product cost Forty percent of fixed manufacturing overhead consists of wages and salaries; the remainder consists of depreciation charges on production equipment and buildings. Production and cost data for the two years are: Units produced Units sold 1 Year 2 20,500 20,500 15,50025,500 Required: 1. Prepare a variable costing contribution format income statement for each year. Heaton Company Variable Costing Income Statement Year 1 Year 2

media%2F148%2F1483bc94-ff2a-4e20-aaf1-1e
media%2Fff2%2Fff26edc1-9c1d-4909-8115-79
Answer #1

Year 1(15500)

Year 2(25,500)

Sales

930,000

$1,530,000

Less:Variable expenses:

Variable COGS sold(6+9+1=16)

248000

408000

Variable selling & adm exp (300700-254200)/15500=$3

46500

76500

Contribution margin

635,500

1,045,500

Fixed expenses:

Fixed manufacturing overhead

369000

369000

Selling & adm expense

254200

254200

Total ficed exp

623200

623200

Net operatimg income

$12,300

$422,300

year 1

year 2

variable costing net income

$12,300

$422,300

fixed overhead deferred

90000

-90,000

(369000/20500)*5000

Absorption costing net operating income

$102,300

332,300

Working notes for the above answer

W.N.1

Sale units

Year-1

=$930,000/60

=15,500 units

Year-2

=1530,000/60

=25,500 units